# Where Did It All Go Wrong? A Hierarchical Look into Multi-Agent Error Attribution

### v1

@@ A.5 Objective Analysis Agent
␣␣␣␣ObjectiveAnalysisAgentthatanalyzesallagentsinaconversationobjectively
␣␣␣␣todetermineerrorattributionwithoutforcedbias.
@@
␣␣␣␣␣␣␣␣InitializetheObjectiveAnalysisAgent.
␣␣␣␣␣␣␣␣Args:
␣␣␣␣␣␣␣␣␣␣␣␣model_id:ThemodelIDtousefortheagent
␣␣␣␣␣␣␣␣␣␣␣␣temperature:Thetemperaturetouseforgeneration
@@
␣␣␣␣␣␣␣␣YouareanObjectiveAnalysisAgentconductinganimpartialinvestigationtodetermineerrorattributioninamulti-agentconversation.
␣␣␣␣␣␣␣␣ANALYSTSPECIALIZATION:{focus_instructions}
␣␣␣␣␣␣␣␣Yourtask:
␣␣␣␣␣␣␣␣1.AnalyzeALLagentsintheconversationobjectively(notjustonespecificagent)
␣␣␣␣␣␣␣␣2.Determinewhichagent(s)mostlikelycausedthefinalwronganswer
␣␣␣␣␣␣␣␣3.Determinewhichstep/turnintheconversationthemistakeoccurred
␣␣␣␣␣␣␣␣4.Provideconfidencescoresandreasoningforyourconclusions
␣␣␣␣␣␣␣␣Youhaveaccesstohierarchicalcontextshowing:
␣␣␣␣␣␣␣␣-Immediateagents:Fulldetails
␣␣␣␣␣␣␣␣-Nearbyagents:Keydecisions
␣␣␣␣␣␣␣␣-Distantagents:Briefsummaries
␣␣␣␣␣␣␣␣Theagentsarenumberedsequentially(Agent1,Agent2,etc.)correspondingtotheirstep/turnindexintheconversation.
␣␣␣␣␣␣␣␣Possibleconclusions:
␣␣␣␣␣␣␣␣-Singleagenterror:Onespecificagentcausedthemistakeataspecificstep
␣␣␣␣␣␣␣␣-Multi-agenterror:Multipleagentscontributedtothemistakeacrossspecificsteps
␣␣␣␣␣␣␣␣OutputyourresponseasvalidJSONwrappedin<json></json>tags:
␣␣␣␣␣␣␣␣<json>
␣␣␣␣␣␣␣␣␣␣"analysis_summary":"Brief overview of your investigation approach and findings",
␣␣␣␣␣␣␣␣␣␣"agent_evaluations":[
␣␣␣␣␣␣␣␣␣␣␣␣␣␣"agent_name":"agent_name",
␣␣␣␣␣␣␣␣␣␣␣␣␣␣"step_index":1,
␣␣␣␣␣␣␣␣␣␣␣␣␣␣"error_likelihood":0.0-1.0,
␣␣␣␣␣␣␣␣␣␣␣␣␣␣"reasoning":"Why this agent may or may not have caused the error",
␣␣␣␣␣␣␣␣␣␣␣␣␣␣"evidence":"Specific evidence supporting your assessment"
␣␣␣␣␣␣␣␣␣␣"primary_conclusion":{{
␣␣␣␣␣␣␣␣␣␣␣␣"type":"single_agent"|"multi_agent",
␣␣␣␣␣␣␣␣␣␣␣␣"attribution":["agent_name(s)"]ornull,
␣␣␣␣␣␣␣␣␣␣␣␣"mistake_step":1,
␣␣␣␣␣␣␣␣␣␣␣␣"confidence":0.0-1.0,
␣␣␣␣␣␣␣␣␣␣␣␣"reasoning":"Detailed explanation of your primary conclusion including which step the error occurred"
␣␣␣␣␣␣␣␣␣␣"alternative_hypotheses":[
␣␣␣␣␣␣␣␣␣␣␣␣␣␣"type":"conclusion_type",
␣␣␣␣␣␣␣␣␣␣␣␣␣␣"attribution":["agent_name(s)"]ornull,
␣␣␣␣␣␣␣␣␣␣␣␣␣␣"mistake_step":1,
␣␣␣␣␣␣␣␣␣␣␣␣␣␣"confidence":0.0-1.0,
␣␣␣␣␣␣␣␣␣␣␣␣␣␣"reasoning":"Alternative explanation"
␣␣␣␣␣␣␣␣</json>
␣␣␣␣␣␣␣␣Bethorough,objective,andconsiderallpossibilitiesincludingthatnosingleagentmaybeclearlyatfault.
␣␣␣␣␣␣␣␣Payspecialattentiontoidentifyingthespecificstep/turnwheretheerroroccurred.
@@
␣␣␣␣␣␣␣␣Analyzetheentireconversationobjectivelytodetermineerrorattribution.
␣␣␣␣␣␣␣␣Args:
␣␣␣␣␣␣␣␣␣␣␣␣conversation_contexts:Listofhierarchicalcontextdictionariesforallagents
␣␣␣␣␣␣␣␣␣␣␣␣ground_truth:Thegroundtruthanswer
␣␣␣␣␣␣␣␣␣␣␣␣final_answer:Thefinalanswergiven
␣␣␣␣␣␣␣␣␣␣␣␣query:Theoriginalquery/question
␣␣␣␣␣␣␣␣Returns:
␣␣␣␣␣␣␣␣␣␣␣␣Dictionarycontainingobjectiveanalysisresults
@@
␣␣␣␣␣␣␣␣OriginalQuery:{query}
␣␣␣␣␣␣␣␣{ground_truth_section}
␣␣␣␣␣␣␣␣FinalAnswer:{final_answer}
␣␣␣␣␣␣␣␣ConversationAnalysis:
␣␣␣␣␣␣␣␣{context_summary}
␣␣␣␣␣␣␣␣Pleaseconductanobjectiveanalysisofthisconversationtodetermineerrorattribution.
␣␣␣␣␣␣␣␣Focusonidentifyingwhichspecificagent(s)causedtheerrorthatledtotheincorrectfinalanswer.
␣␣␣␣␣␣␣␣OutputyouranalysisintheJSONformatspecifiedinyourinstructions.
@@
␣␣␣␣␣␣␣␣Createacomprehensivesummaryoftheconversationforobjectiveanalysis.
␣␣␣␣␣␣␣␣Args:
␣␣␣␣␣␣␣␣␣␣␣␣conversation_contexts:Listofhierarchicalcontextdictionaries
␣␣␣␣␣␣␣␣Returns:
␣␣␣␣␣␣␣␣␣␣␣␣Formattedconversationsummary
@@
␣␣␣␣␣␣␣␣Getspecializedinstructionsbasedonanalystfocus.
␣␣␣␣␣␣␣␣Args:
␣␣␣␣␣␣␣␣␣␣␣␣analyst_focus:Thetypeofanalystfocus
␣␣␣␣␣␣␣␣Returns:
␣␣␣␣␣␣␣␣␣␣␣␣Specializedinstructionsstring
@@ A.6 Judge Agent
␣␣␣␣FinalJudgeAgentthatweighscompetingargumentsfrommultipleContext-AwareStepAgents
␣␣␣␣todeterminethetrueerrorattribution.
@@
␣␣␣␣␣␣␣␣InitializetheFinalJudgeAgent.
␣␣␣␣␣␣␣␣Args:
␣␣␣␣␣␣␣␣␣␣␣␣model_id:ThemodelIDtousefortheagent
␣␣␣␣␣␣␣␣␣␣␣␣temperature:Thetemperaturetouseforgeneration(lowerformoredeterministicoutput)
@@
␣␣␣␣␣␣␣␣YouareaFinalJudgeAgentthatweighscompetingargumentsfrommultipleContext-AwareStepAgentstodeterminethetrueerrorattribution.
␣␣␣␣␣␣␣␣Yourtask:EachContext-AwareStepAgenthasarguedwhyTHEIRagentcausedtheerror.Reviewallargumentsanddeterminewhichoneismostconvincingbasedonevidenceandreasoning.
␣␣␣␣␣␣␣␣TheargumentsfromeachContext-AwareStepAgentareprovidedinastructuredtextformatwiththesesections:
␣␣␣␣␣␣␣␣-Purpose:Thepurposeoftheagentstep
␣␣␣␣␣␣␣␣-AssumptionsandInformation:Whattheagentwasgiven
␣␣␣␣␣␣␣␣-Errors:Whattheagentdidwrong(ifanything)
␣␣␣␣␣␣␣␣-Evidence:Evidencesupportingtheerrorattribution
␣␣␣␣␣␣␣␣OutputyourresponseasavalidJSONobjectwrappedin<json></json>XMLtags.TheJSONshouldhavethefollowingstructure:
␣␣␣␣␣␣␣␣<json>
␣␣␣␣␣␣␣␣␣␣"mistake_agent":"agent_name",
␣␣␣␣␣␣␣␣␣␣"mistake_step":"step_number",
␣␣␣␣␣␣␣␣␣␣"mistake_reason":"explanation of why this agent/step caused the wrong final answer, based on the most convincing argument"
␣␣␣␣␣␣␣␣</json>
␣␣␣␣␣␣␣␣IMPORTANTRULES:
␣␣␣␣␣␣␣␣1.YourresponseMUSTbeavalid,parsableJSONobjectwrappedin<json></json>tags.Donotincludeanytextoutsidethesetags.
␣␣␣␣␣␣␣␣2.Focusontheagentsthatareactivelymakingdecisionsorprovidinginformation.
␣␣␣␣␣␣␣␣Bethoroughinyouranalysis.Considerthestrengthofevidence,thelogicalconnectionbetweentheerrorandthefinalwronganswer,andthecausalrelationship.
␣␣␣␣␣␣␣␣Workbackwardstoseewherethelogicdivergedandtheerrorhappened.
@@
␣␣␣␣␣␣␣␣Judgethecompetingargumentsanddeterminethetrueerrorattribution.
␣␣␣␣␣␣␣␣Args:
␣␣␣␣␣␣␣␣␣␣␣␣agent_arguments:ListofargumentsfromContext-AwareStepAgents
␣␣␣␣␣␣␣␣␣␣␣␣ground_truth:Thegroundtruthanswer
␣␣␣␣␣␣␣␣␣␣␣␣final_answer:Thefinalanswergiven
␣␣␣␣␣␣␣␣␣␣␣␣query:Theoriginalquery/question
␣␣␣␣␣␣␣␣Returns:
␣␣␣␣␣␣␣␣␣␣␣␣Finalerrorattributionasadictionary
@@
␣␣␣␣␣␣␣␣OriginalQuery:{query}
␣␣␣␣␣␣␣␣GroundTruth:{ground_truth}
␣␣␣␣␣␣␣␣FinalAnswer:{final_answer}
␣␣␣␣␣␣␣␣AllAgentArguments:{agent_arguments_str}
␣␣␣␣␣␣␣␣PleasereviewalltheargumentsfromtheContext-AwareStepAgentsanddeterminewhichoneismostconvincing.
␣␣␣␣␣␣␣␣OutputyourdecisionasavalidJSONobjectwrappedin<json></json>XMLtagsasspecifiedinyourinstructions.
␣␣␣␣␣␣␣␣IMPORTANT:YourresponseMUSTbeavalid,parsableJSONobjectwrappedin<json></json>tags.Donotincludeanytextoutsidethesetags.
@@ A.7 Consensus Voting
␣␣␣␣ConsensusVotingAgentthataggregatesmultipleobjectiveanalyses
␣␣␣␣todeterminefinalerrorattributionthroughvoting.
@@
␣␣␣␣␣␣␣␣InitializetheConsensusVotingAgent.
␣␣␣␣␣␣␣␣Args:
␣␣␣␣␣␣␣␣␣␣␣␣min_confidence_threshold:Minimumconfidencethresholdtoconsideraconclusion
@@
␣␣␣␣␣␣␣␣Aggregatemultipleobjectiveanalysesthroughconsensusvoting.
␣␣␣␣␣␣␣␣Args:
␣␣␣␣␣␣␣␣␣␣␣␣objective_analyses:Listofobjectiveanalysisresults
␣␣␣␣␣␣␣␣␣␣␣␣ground_truth:Thegroundtruthanswer
␣␣␣␣␣␣␣␣␣␣␣␣final_answer:Thefinalanswergiven
␣␣␣␣␣␣␣␣␣␣␣␣query:Theoriginalquery/question
␣␣␣␣␣␣␣␣Returns:
␣␣␣␣␣␣␣␣␣␣␣␣Dictionarycontainingconsensusattributionresults
@@
␣␣␣␣␣␣␣␣Performconsensusvotingontheanalyses.
␣␣␣␣␣␣␣␣Args:
␣␣␣␣␣␣␣␣␣␣␣␣primary_conclusions:Listofprimaryconclusionsfromanalysts
␣␣␣␣␣␣␣␣␣␣␣␣agent_evaluations:Dictionaryofagentevaluationsbyagentname
␣␣␣␣␣␣␣␣␣␣␣␣alternative_hypotheses:Listofalternativehypotheses
␣␣␣␣␣␣␣␣Returns:
␣␣␣␣␣␣␣␣␣␣␣␣Consensusvotingresults
@@
␣␣␣␣␣␣␣␣Analyzedisagreementsbetweenanalysts.
␣␣␣␣␣␣␣␣Args:
␣␣␣␣␣␣␣␣␣␣␣␣conclusion_votes:Dictionaryofconclusionvotes
␣␣␣␣␣␣␣␣Returns:
␣␣␣␣␣␣␣␣␣␣␣␣Disagreementanalysis
@@
␣␣␣␣␣␣␣␣Synthesizereasoningfrommultipleanalystvotes.
␣␣␣␣␣␣␣␣Args:
␣␣␣␣␣␣␣␣␣␣␣␣best_conclusion_info:Informationaboutthebestconclusion
␣␣␣␣␣␣␣␣Returns:
␣␣␣␣␣␣␣␣␣␣␣␣Synthesizedreasoning
@@
␣␣␣␣␣␣␣␣Createanemptyresultwhennoanalysesareprovided.
␣␣␣␣␣␣␣␣Returns:
␣␣␣␣␣␣␣␣␣␣␣␣Emptyconsensusresult

### v2

@@ A.5 Objective Analysis Agent
Objective Analysis Agent that analyzes all agents in a conversation objectively
to determine error attribution without forced bias.
@@
Initialize the Objective Analysis Agent.
Args:
model_id: The model ID to use for the agent
temperature: The temperature to use for generation
@@
You are an Objective Analysis Agent conducting an impartial investigation to determine error attribution in a multi-agent conversation.
ANALYST SPECIALIZATION: {focus_instructions}
Your task:
1. Analyze ALL agents in the conversation objectively (not just one specific agent)
2. Determine which agent(s) most likely caused the final wrong answer
3. Determine which step/turn in the conversation the mistake occurred
4. Provide confidence scores and reasoning for your conclusions
You have access to hierarchical context showing:
- Immediate agents: Full details
- Nearby agents: Key decisions
- Distant agents: Brief summaries
The agents are numbered sequentially (Agent 1, Agent 2, etc.) corresponding to their step/turn index in the conversation.
Possible conclusions:
- Single agent error: One specific agent caused the mistake at a specific step
- Multi-agent error: Multiple agents contributed to the mistake across specific steps
Output your response as valid JSON wrapped in <json></json> tags:
<json>
"analysis_summary": "Brief overview of your investigation approach and findings",
"agent_evaluations": [
"agent_name": "agent_name",
"step_index": 1,
"error_likelihood": 0.0-1.0,
"reasoning": "Why this agent may or may not have caused the error",
"evidence": "Specific evidence supporting your assessment"
"primary_conclusion": {{
"type": "single_agent" | "multi_agent",
"attribution": ["agent_name(s)"] or null,
"mistake_step": 1,
"confidence": 0.0-1.0,
"reasoning": "Detailed explanation of your primary conclusion including which step the error occurred"
"alternative_hypotheses": [
"type": "conclusion_type",
"attribution": ["agent_name(s)"] or null,
"mistake_step": 1,
"confidence": 0.0-1.0,
"reasoning": "Alternative explanation"
</json>
Be thorough, objective, and consider all possibilities including that no single agent may be clearly at fault.
Pay special attention to identifying the specific step/turn where the error occurred.
@@ 4. Provide confidence scores and reasoning for your conclusions
Analyze the entire conversation objectively to determine error attribution.
Args:
conversation_contexts: List of hierarchical context dictionaries for all agents
ground_truth: The ground truth answer
final_answer: The final answer given
query: The original query/question
Returns:
Dictionary containing objective analysis results
@@
Original Query: {query}
{ground_truth_section}
Final Answer: {final_answer}
Conversation Analysis:
{context_summary}
Please conduct an objective analysis of this conversation to determine error attribution.
Focus on identifying which specific agent(s) caused the error that led to the incorrect final answer.
Output your analysis in the JSON format specified in your instructions.
@@
Create a comprehensive summary of the conversation for objective analysis.
Args:
conversation_contexts: List of hierarchical context dictionaries
Returns:
Formatted conversation summary
@@
Get specialized instructions based on analyst focus.
Args:
analyst_focus: The type of analyst focus
Returns:
Specialized instructions string
@@ A.6 Judge Agent
Final Judge Agent that weighs competing arguments from multiple Context-Aware Step Agents
to determine the true error attribution.
@@
Initialize the Final Judge Agent.
Args:
model_id: The model ID to use for the agent
temperature: The temperature to use for generation (lower for more deterministic output)
@@
You are a Final Judge Agent that weighs competing arguments from multiple Context-Aware Step Agents to determine the true error attribution.
Your task: Each Context-Aware Step Agent has argued why THEIR agent caused the error. Review all arguments and determine which one is most convincing based on evidence and reasoning.
The arguments from each Context-Aware Step Agent are provided in a structured text format with these sections:
- Purpose: The purpose of the agent step
- Assumptions and Information: What the agent was given
- Errors: What the agent did wrong (if anything)
- Evidence: Evidence supporting the error attribution
Output your response as a valid JSON object wrapped in <json></json> XML tags. The JSON should have the following structure:
<json>
"mistake_agent": "agent_name",
"mistake_step": "step_number",
"mistake_reason": "explanation of why this agent/step caused the wrong final answer, based on the most convincing argument"
</json>
IMPORTANT RULES:
1. Your response MUST be a valid, parsable JSON object wrapped in <json></json> tags. Do not include any text outside these tags.
2. Focus on the agents that are actively making decisions or providing information.
Be thorough in your analysis. Consider the strength of evidence, the logical connection between the error and the final wrong answer, and the causal relationship.
Work backwards to see where the logic diverged and the error happened.
@@
Judge the competing arguments and determine the true error attribution.
Args:
agent_arguments: List of arguments from Context-Aware Step Agents
ground_truth: The ground truth answer
final_answer: The final answer given
query: The original query/question
Returns:
Final error attribution as a dictionary
@@
Original Query: {query}
Ground Truth: {ground_truth}
Final Answer: {final_answer}
All Agent Arguments: {agent_arguments_str}
Please review all the arguments from the Context-Aware Step Agents and determine which one is most convincing.
Output your decision as a valid JSON object wrapped in <json></json> XML tags as specified in your instructions.
IMPORTANT: Your response MUST be a valid, parsable JSON object wrapped in <json></json> tags. Do not include any text outside these tags.
@@ A.7 Consensus Voting
Consensus Voting Agent that aggregates multiple objective analyses
to determine final error attribution through voting.
@@
Initialize the Consensus Voting Agent.
Args:
min_confidence_threshold: Minimum confidence threshold to consider a conclusion
@@
Aggregate multiple objective analyses through consensus voting.
Args:
objective_analyses: List of objective analysis results
ground_truth: The ground truth answer
final_answer: The final answer given
query: The original query/question
Returns:
Dictionary containing consensus attribution results
@@
Perform consensus voting on the analyses.
Args:
primary_conclusions: List of primary conclusions from analysts
agent_evaluations: Dictionary of agent evaluations by agent name
alternative_hypotheses: List of alternative hypotheses
Returns:
Consensus voting results
@@
Analyze disagreements between analysts.
Args:
conclusion_votes: Dictionary of conclusion votes
Returns:
Disagreement analysis
@@
Synthesize reasoning from multiple analyst votes.
Args:
best_conclusion_info: Information about the best conclusion
Returns:
Synthesized reasoning
@@
Create an empty result when no analyses are provided.
Returns:
Empty consensus result